The value of sin^-1 ( sin 10 ), isa) 10b) 10-3πc) 3π-10d) none of these

Grade:

1 Answers

Arun
5 years ago
Know that the range of the sin⁻¹ function is [-π/2, π/2]. You need to find the value in that interval having the same sine as 10. 

sin(10) = sin(π - 10) = sin(3π - 10) 
-π/2 ≤ 3π - 10 ≤ π/2 
sin⁻¹[sin(10)] = 3π - 10